OpenStudy (anonymous):

Check my answers pls Estimate by first rounding each number to the nearest 1/2 . 3 4/11 x 2 1/10 6 <---- Amit wants to estimate 1 1/9 + 3 6/7 . To estimate, he simplifies the expression 1 + 4 . Is Amit’s estimate a close estimate, an underestimate, or an overestimate? Underestimate <----- Joy has a piece of ribbon that is 6 7/8 in. long. She purchases a piece that is 35 7/8 in. long. Which expression should Joy use to estimate how many total inches of ribbon she has altogether? 36 + 7 <---- I only have 1 more

OpenStudy (anonymous):

Jacob’s recipe calls for 4 7/8 cups of raisins. He has 6/11 of the amount that he needs. About how many cups of raisins does Jacob have? Estimate by first rounding each number to the nearest 1/2 .

OpenStudy (anonymous):

My answer is 5 1/2 cups

OpenStudy (anonymous):

I don't know but i'd say you did good xD

OpenStudy (anonymous):

oh ok haha I'll ask someone else though to make sure

OpenStudy (anonymous):

@Owlcoffee @Directrix @jim_thompson5910

jimthompson5910 (jim_thompson5910):

how would you round `3 4/11`

jimthompson5910 (jim_thompson5910):

keep in mind that 4/11 = 0.36363636...

jimthompson5910 (jim_thompson5910):

oh nvm, I'm thinking of something else

jimthompson5910 (jim_thompson5910):

3 4/11 rounds to 3 2 1/10 rounds to 2 so we have 3*2 = 6 I agree with the first one

OpenStudy (anonymous):

sorry I was gone for a sec, are u still checking?

jimthompson5910 (jim_thompson5910):

yeah

jimthompson5910 (jim_thompson5910):

the second one is hard to tell unless you use a calculator. I'm not sure what your teacher had in mind here

OpenStudy (anonymous):

To tell you the truth, I looked up the answer for that one. I didn't know how to do it

jimthompson5910 (jim_thompson5910):

anyways, I'm onto the last one what does 4 7/8 round to?

OpenStudy (anonymous):

1?

jimthompson5910 (jim_thompson5910):

4&7/8 not just 7/8

OpenStudy (anonymous):

oh, 5?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

6/11 = 0.54545454... what does 6/11 round to?

OpenStudy (anonymous):

1/2

jimthompson5910 (jim_thompson5910):

so you'll be multiplying 5 with 1/2 (the answer isn't 5&1/2)

OpenStudy (anonymous):

5/2?

jimthompson5910 (jim_thompson5910):

or 2&1/2 correct

OpenStudy (anonymous):

@jim_thompson5910 So is everything right except the last one? And how is it 2 1/2?

jimthompson5910 (jim_thompson5910):

5/2 = (4+1)/2 = (4/2) + (1/2) = (2) + (1/2) = 2&1/2

OpenStudy (anonymous):

ok thank you

jimthompson5910 (jim_thompson5910):

no problem